Let A and B be any two events. If P(A) = 0.2, P(B) = 0.8 and P(A and B) = 0.6, Which probability is possible

Select one:

a. both

b. P(B|A)

c. none

d. P(A|B)
Answer #1

P(A) = 0.2
P(B) = 0.8
P(A and B) = 0.6


P(A|B) = P(A and B)/P(B) = 0.6/0.8 = 0.75

P(B|A) = P(A and B)/P(A) = 0.6/0.2 = 3

As P(B|A) > 1, it is not possible.
P(A|B) = 0.75 < 1 which is possible.

Correct option: D) P(A|B)